OpenStudy (anonymous):

If you roll two number cubes, what is the probability of their sum being seven or higher? 7/12 5/12 7/5 5/7

OpenStudy (anonymous):

your possible results: 1+1=2 1+2=3 1+3=4 1+4=5 1+5=6 1+6=7 2+1=3 2+2=4 2+3=5 2+4=6 2+5=7 2+6=8 3+1=4 3+2=5 3+3=6 3+4=7 3+5=8 3+6=9 4+1=5 4+2=6 4+3=7 4+4=8 4+5=9 4+6=10 5+1=6 5+2=7 5+3=8 5+4=9 5+5=10 5+6=11 6+1=7 6+2=8 6+3=9 6+4=10 6+5=11 6+6=12

OpenStudy (anonymous):

these are 36 (6^2) possible outcomes, now only count how many of them are greater or equal to 7

OpenStudy (anonymous):

sooo 7/12?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

21/36